An erlenmeyer flask contains 50.0 mL of an aqueous solution of 0.425 M acetic acid (CH3COOH). Given that acetic acid has a Ka = 1.75 x10^(-5), answer the questions that follow.


(A) Calculate the pH of the resulting solution when the original solution is diluted to 250.0 mL via addition of distilled water. Write answer in THREE SIGNIFICANT FIGURES and complete solutions.


1
Expert's answer
2021-04-12T03:46:10-0400

Given data-

Ka of acetic acid=1.75×10^-5

Concentration of acetic acid= 0.425 M

Given equation -

CH3COOH=CH3COO- + H+

Calculate H+ using Ka equation -

Ka =[H+]2 / acid

1.75×10^-5=[H+]2 /. 425

[H+]2 =1.75×10^-5 ×. 425=7.4375×10^-6

[H+] = 2.72×10^-3 M

We know,

PH =-log[H+] = -log[2.72×10^-3]

PH =2.56

Hence the PH of the acetic acid is 2.56